OpenStudy (anonymous):

NEGATIVE 5 RAISED TO ONE FIFTH POWER

OpenStudy (zarkon):

\[\sqrt[5]{5}e^{\frac{2\pi i k}{5}-\frac{\pi}{5}i}\] \[\sqrt[5]{5}e^{\frac{\pi i (2k-1)}{5}}\]

OpenStudy (zarkon):

\[k\in\mathbb{Z}\]